OpenStudy (anonymous):

Let f(x) = sqrt(x+2). Calculate the difference quotient f(14+h)-f(14)/h for: h= .1 The slope of the tangent line to the graph of f(x) at x=14 is m= The equation of the tangent line to the curve at the point (14, 4 ) is y=

myininaya (myininaya):

\[\frac{f(x+h)-f(x)}{h}=\frac{\sqrt{14+h+2}-\sqrt{x+2}}{h} \]